Two students are experimenting on falling bodies. From a window 15 m above the ground, one student drops a ball and, at the same time, the other student throws another ball vertically up with an initial speed of 9.5 m/s. How much time would elapse between the balls striking the ground?




1
Expert's answer
2021-09-07T09:50:52-0400
"t_1=\sqrt{\frac{2(15)}{9.8}}=1.75\ s\\
-15=9.5t_2-0.5(9.8)t_2^2\\t_2=2.97\ s\\"